# Federated Learning for Smart Grid: A Survey on Applications and Potential Vulnerabilities

### v2

@@ 4.2. Vulnerabilities in Transmission and Distribution Stages
SG technology revolutionizes traditional power grids by incorporating advanced ICTs into the existing energy infrastructure. This integration enables a more efficient, reliable, and flexible system for electricity transmission and distribution (hamidi2010smart). The transmission component of the SG (bayoumi2015power) focuses on the high-voltage transport of electricity over long distances. In SGs, transmission systems benefit from advanced technologies such as Flexible AC Transmission Systems (FACTS) and High-Voltage Direct Current (HVDC) power transmission, which enhance the controllability, stability, and power transfer capability of AC transmission systems. However, transmission systems are vulnerable to various types of attacks. FDIAs are particularly concerning, as they can manipulate sensor measurements and mislead grid operators.
SG distribution networks (konstantelos2016strategic) are responsible for delivering electricity to end-users at lower voltages. These networks are undergoing significant changes to accommodate the increasing integration of DERs and renewable energy sources. Smart distribution systems enable bi-directional power flow, allowing consumers to also become producers of electricity. However, distribution systems face threats such as energy theft through meter tampering, privacy breaches of consumer data, and coordinated attacks on demand response systems (ebtia2024spatial).
4.2.1. State Estimation
@@
AC State Estimation is a crucial process in SG systems, designed to determine the most likely steady-state operation of the power system based on field measurements. It serves as a fundamental data processing tool responsible for supporting and increasing system visibility while filtering errors that may appear in real-time measurements, system topology, and parameters (acurio2023state).
@@
However, despite its significance, AC State Estimation is vulnerable to cyber threats, particularly FDIAs. These attacks can manipulate measurement data, introduce errors in state estimation, which can impact bus power flow calculations and the economic dispatch of the power system (huang2022false). Within the FL framework, Lin et al. (lin2022incentive) propose an edge-based FL framework for detecting FDIAs in SG state estimation. This approach leverages local system operating data to predict and assess system stability in a distributed manner. By utilizing edge devices for local model training, the framework reduces the computational burden on central servers and enhances privacy protection. A key innovation in this work is the incorporation of an incentive mechanism designed to encourage high-quality data sharing and model training. This mechanism addresses the challenge of motivating participants to actively contribute to the FL process, which is essential for improving the overall detection accuracy of the system.
4.2.2. Events Classification and Anomaly Detection
Grid events classification involves identifying and categorizing various disturbances or anomalies that occur within power systems. This process is crucial for maintaining the reliability, stability, and security of electrical grids. By accurately detecting and classifying these events, power system operators can quickly respond to potential issues and mitigate their impacts on the grid operations.
Deep learning-based models (niazazari2020attack), which have gained popularity for grid event classification, are particularly vulnerable to adversarial attacks. While highly effective under normal conditions, these models can be deceived by carefully crafted perturbations in the input data. The vulnerability arises from the complex nature of these models and their sensitivity to small changes in input features. Within the FL framework, Mohammadabadi et al. (mohammadabadi2023towards) propose a federated XGBoost model to address data privacy concerns and enable collaborative learning. One key challenge in implementing FL for phasor measurement unit (PMU) data is the varying numbers of PMUs across different entities, resulting in diverse input structures for the event classification model. To overcome this, the proposed model constructs a common set of event features that can be generated by each entity, regardless of how many PMUs they possess. This study also considers two types of adversarial attacks on images using the Fast Gradient Sign Method (FGSM) and an iterative variant. In both approches, perturbations are applied to clean images to generate adversarial examples. For untargeted attacks, FGSM updates the adversarial image as follows:
(7) $x_{\text{adv}}=x+\epsilon\cdot\text{sign}(\nabla_{x}L(x,y)),$
where $\epsilon$ controls the perturbation size, $L(x,y)$ is the loss, and $\nabla_{x}$ represents the gradient. For targeted attacks, an iterative method is used:
(8) $x_{\text{adv}}=x+\Delta,$
where $\Delta_{i+1}=\Delta_{i}-\alpha\cdot\text{sign}(\nabla_{x}L(x,y))$ for each iteration $i$, with constraints $-\epsilon\leq\Delta_{i+1}\leq\epsilon$. These perturbations are visually subtle but malicious, creating adversarial images that fool machine learning models. To mitigate theses attacks, the paper employs data processing techniques and a LSTM network. To further strengthen privacy protection in the FL process, the paper incorporates an additive homomorphic encryption scheme. This encryption method secures the communication during the training process, eliminating the need for a trusted server and enhancing overall data protection.
Jithish et al (jithish2023distributed) propose a FL-based solution that enables the training of machine learning models across multiple distributed devices without compromising user privacy or relying on extensive centralized computing resources. This approach not only preserves data confidentiality but also enables real-time anomaly detection, which is crucial for identifying and responding to various grid issues such as unusual power consumption patterns, infrastructure faults, and potential attacks. In this system, a global model is downloaded from the server to smart meters for on-device training. After local training, only the updated local model parameters are sent to the server to improve the global model, thus ensuring user privacy. To further protect the model parameter updates from potential adversaries, the SSL/TLS protocol is employed. adding an extra layer of security to guarantee the integrity and confidentiality of the transmitted data.
4.2.3. Voltage Control
@@
Voltage regulation is a critical process in electrical power systems, designed to maintain voltage levels within acceptable limits across the network. It involves controlling and adjusting the voltage to ensure stable and reliable power delivery to consumers (xie2021review). However, these networks are susceptible to data-related attacks, which can lead to voltage instability. Attackers may exploit vulnerabilities in communication channels to inject false data or disrupt the flow of information critical for effective voltage control.
@@
In the FL-based SG system, Liu et al. (liu2022federated) combine FL with reinforcement learning techniques to develop a privacy-preserving and efficient control strategy. In this federated reinforcement learning framework, grid-connected interface inverters of renewable energy sources are modeled as deep neural network-based agents that interact with their local environment to learn optimal voltage-var control strategies. The process operates in a two-stage process: local training using reinforcement learning algorithms, followed by periodic global aggregation of model updates. This approach offers several advantages, including enhanced data privacy, reduced communication costs, improved scalability, and adaptability to changing network conditions.
4.3. Vulnerabilities in Consumption Stage
@@
The consumption part of a SG represents the end-user side of the energy distribution system (kaur2020energy), encompassing homes, businesses, industries, and more. It integrates advanced technologies like AMI with smart meters for real-time usage monitoring and demand-side management to optimize energy consumption patterns. This system leverages IoT devices and real-time analytics to enhance energy efficiency and provide consumers with detailed insights into their usage. Additionally, sophisticated forecasting techniques and machine learning algorithms are employed to predict and manage energy loads effectively, ensuring a balance between supply and demand. However, smart meters and IoT devices, which are integral to these systems, can be vulnerable to attacks (liu2021vulnerability), potentially allowing malicious customers to manipulate energy consumption data or disrupt power supply.
4.3.1. Load Forecasting
@@
Smart meter measurements, while essential for accurate demand forecasting, raise significant privacy and data breach concerns for consumers. The fine-grained load data collected by smart meters can potentially reveal sensitive information about residential users’ daily routines and behaviors, posing risks to property security (husnoo2023secure). In the FL framework, Husnoo et al. (husnoo2023secure) propose a secure FL approach designed for residential short-term load forecasting. This approach addresses the critical balance between accurate demand prediction and protecting consumer privacy in smart meter data. FL allows collaborative model training without exposing raw data, as models are trained locally, and only parameters are shared with a central server. To enhance security, the framework incorporates advanced techniques like gradient quantization using SignSGD (bernstein2018signsgd) and differential privacy (dwork2006differential). Additionally, it employs hierarchical and peer-to-peer strategies to improve both privacy and accuracy. The system tackles challenges such as non-IID data distribution and potential attacks through methods like deep federated adaptation and secure aggregation. Experimental results demonstrate that this approach achieves high forecasting accuracy while preserving user data privacy, with significant improvements in prediction accuracy compared to traditional methods.
4.3.2. Energy Theft Detection
@@
Energy theft detection (salinas2015privacy) in SGs refers to the process of identifying and preventing unauthorized consumption of electricity, an illegal activity that causes significant economic losses to utility companies. SGs integrate digital technologies with traditional power systems, enabling advanced energy management and monitoring capabilities. These systems utilize AMI and smart meters to collect, measure, and analyze energy usage data from customers at regular intervals. However, the collection and analysis of consumer energy data create vulnerabilities, potentially allowing attackers to compromise smart meters.
@@
Within the FL framework, Wen et al. (wen2021feddetect) introduce FedDetect, a novel privacy-preserving FL framework designed for energy theft detection in SGs. FedDetect leverages FL to train machine learning models on distributed devices, aggregating local model updates instead of raw data to protect user privacy. Ashraf et al. (ashraf2022feddp) propose FedDP, a privacy-protecting theft detection scheme for SGs using FL. This framework addresses the challenge of utilizing consumer energy data for theft detection while maintaining privacy protection. Zafar et al. (zafar2023step) develop a FL-assisted hybrid deep learning model for electricity theft detection in SGs. Their model combines FL techniques with deep learning to enhance security and reliability of electricity theft detection. Li et al. (li2024fuse) present FUSE, a novel electricity theft detection framework that integrates FL with U-shape split learning. FUSE aims to address the challenges of data privacy and security in SG systems while effectively detecting electricity theft.
5. Discussions
@@
5.1. SOTA FL Attack
5.1.1. Byzantine Attack in FL
@@
Byzantine attacks (aka model poisoning attacks) in FL involve malicious clients sending incorrect or misleading model updates to the central server, aiming to disrupt the training process and degrade model performance. These malicious local model updates are typically intentionally modified by those malicious clients controlled by the attacker. We introduce several classical and commonly discussed Byzantine attacks as follows.
Random attack. To disturb the convergence of the global model, the malicious clients send randomized updates that follow a Gaussian distribution $N(\mu,\sigma^{2}\textbf{I}_{d})$ where $d$ is the dimension of the local model update. In particular, $\mu$ and $\sigma$ should be carefully chosen to make the malicious model updates have a similar distribution to benign ones.
@@
Noise attack. The malicious clients perturb their local model updates by adding Gaussian noise sampled from $N(\mu,\sigma^{2}\textbf{I}_{d})$. Here, $\mu$ and $\sigma$ should also be carefully chosen as a negligible noise (small $\mu$ and $\sigma$) cannot disrupt the convergence of the global model while a large noise (large $\mu$ and $\sigma$) makes malicious model updates significantly differ from benign ones.
Sign-flip attack. Instead of manipulating the magnitude of model updates, the malicious clients change the sign of their model updates to disturb the convergence of the global model. The Sign-flip attack is achieved by flipping the element-wise sign in the model updates.
Min-Max/Min-Sum attack (dnc). The creation of malicious model updates involves two steps. First, the attacker generates a malicious update by perturbing the average of all benign updates. In the second step, specifically for the Min-Max attack, the attacker optimizes this malicious update to ensure that its maximum Euclidean distance from any benign update is no greater than the maximum distance between any two benign updates. As for Min-Sum attack, the malicious update is optimized to ensure that the sum of its distances with each benign update is upper-bounded by the maximum total distance of a benign update among other benign updates. The Min-Max and Min-Sum attacks are powerful since malicious clients try to search for an optimal perturbation such that it has maximum impact.
AGR-tailored Trimmed-mean attack (dnc). In (dnc), the authors proposed several attack methods specific to some well-known defense methods. We take the AGR-tailored Trimmed-mean attack as a example. The AGR-tailored Trimmed-mean is designed to attack the defense method Trimmed-mean proposed in (trmean) by maximizing the Euclidean distance between the aggregated result of simple average and Trmean. This straightforward and intuitive method is effective, but the malicious model updates produced by the AGR-tailored Trimmed-mean attack differ significantly from benign ones, making them susceptible to easy detection by defense methods.
Lie attack (lie). In A-little-is-enough attack, the malicious clients apply slight perturbation to their local benign updates, making it hard to be detected by the defense methods. Specifically, the malicious clients calculate the element-wise mean $\mu_{j}$ and standard error $\sigma_{j}$ of all updates and generate the element of malicious updates by $(\beta_{i})_{j}=\mu_{j}-z\times\sigma_{j}$, where $\beta$ is the malicious model updates and $j\in[d]$. Here, $z$ is a scaling factor decided by the number of malicious clients and the total number of clients.
ByzMean attack (signguard). The ByzMean attack manipulates the mean of updates to be an arbitrarily malicious update. Specifically, it divides the malicious clients into two groups, with $m_{1}$ and $m_{2}$ clients respectively. The first group uses any existing attack methods (e.g., Lie attack) to generate their malicious updates. The second group creates their malicious updates to ensure that the overall average of all updates matches the average of the malicious updates from the first group.
5.1.2. Backdoor Attack in FL
Backdoor attacks (aka data poisoning attacks) in FL aim to maintain the global model’s performance on clean inputs while causing it to make incorrect predictions on inputs containing a specific predefined feature (also called trigger). Since backdoor attacks (lyu2024task; lyu2023attention) do not significantly affect the main task accuracy, the malicious local model updates are similar to benign ones, making anomaly detection more challenging in FL. We introduce several classical and commonly discussed backdoor attacks as follows.
Badnet attack (badnet). Badnet is a simple and effective backdoor attack where the attacker injects the same trigger into the training data of compromised clients. The ground truth label of these training data will also be changed to another backdoor class.
Distributed backdoor attack (dba). In the distributed backdoor attack (DBA), the adversary decomposes the backdoor trigger into several small triggers and applies them to different compromised clients rather than using the same trigger for all malicious clients (Badnet attack). This makes the attack harder to detect, as each client’s updates appear less suspicious.
Scaling attack (scaling). Based on Badnet attack, a scaling attack involves malicious clients scaling their model updates in a way that scales up the impact of their backdoor-triggered updates without affecting the clean data performance. The scaling factor ensures that the backdoor pattern has a stronger influence on the global model.
PGD attack (PGD). The Projected Gradient Descent (PGD) attack is adapted from adversarial machine learning. It iteratively perturbs the model updates of compromised clients within a certain bound to embed the backdoor. This method is effective at creating small but impactful changes to the model parameters.
Neurotoxin attack (Neurotoxin). Neurotoxin explores a persistent attack method designed for scenarios where attackers can only participate in a limited number of rounds. Specifically, in the Neurotoxin attack, the malicious clients project their model updates into a subspace that excludes the global coordinates. This approach ensures that the projected updates from malicious clients are primarily embedded in coordinates with minimal perturbation from benign updates, maintaining the backdoor effect even after the attack ceases.
5.1.3. Inference Attack in FL
@@
Membership Inference Attack (MIA) (membership1; membership2) is a major privacy attack that seeks to determine whether a specific data record was part of the training dataset used to build a machine learning model. In the context of FL, MIA poses a significant threat to the privacy of participating clients and their data. Various entities, including malicious clients, servers, or external attackers, can execute MIA in FL. These adversaries exploit the model’s natural tendency to overfit on training data, leading to unintended information leakage. Approaches to conducting MIA in FL include attack models (membership3) and gradient-based MIAs, which exploit the gradients exchanged during the FL process. Passive MIA techniques also exist, where adversaries infer membership information by merely observing communication between clients and the server. Successful MIAs compromise the privacy of individual users, a particularly serious issue in sensitive domains such as healthcare or finance, where data confidentiality is paramount. Beyond identifying membership, MIA can also reveal attributes or properties of the training data, further threatening user privacy. Defending against MIA in FL presents unique challenges due to the distributed nature of the learning process, with traditional privacy-preserving techniques either being inapplicable or significantly impacting model utility. As a result, researchers are actively developing novel defense mechanisms to improve FL’s resilience to MIA while preserving model performance.
5.2. SOTA FL Defense
To defend against the attacks mentioned above, the server in FL usually adopts a defense method. Existing defense methods can be generally divided into three categories: detection-filtering methods (krum; bulyan; xu2022byzantine; LASA; signguard; dnc; fltrust), impact-reduction methods(trmean; bulyan; Foolsgold; lockdown; geomed; sparsefed; zhang2023byzantine), and differential privacy-based methods.
5.2.1. Detection-Filtering Methods

### v3

@@ 4.1.2. Vulnerabilities in Transmission and Distribution Stages
SG technology revolutionizes traditional power grids by incorporating advanced ICTs into the existing energy infrastructure. This integration enables a more efficient, reliable, and flexible system for electricity transmission and distribution (Hamidi et al., 2010). The transmission component of the SG (Bayoumi, 2015) focuses on the high-voltage transport of electricity over long distances. In SGs, transmission systems benefit from advanced technologies such as Flexible AC Transmission Systems (FACTS) and High-Voltage Direct Current (HVDC) power transmission, which enhance the controllability, stability, and power transfer capability of AC transmission systems. However, transmission systems are vulnerable to various types of attacks. FDIAs are particularly concerning, as they can manipulate sensor measurements and mislead grid operators. SG distribution networks (Konstantelos et al., 2016) are responsible for delivering electricity to end-users at lower voltages. These networks are undergoing significant changes to accommodate the increasing integration of DERs and renewable energy sources. Smart distribution systems enable bi-directional power flow, allowing consumers to also become producers of electricity. However, distribution systems face threats such as energy theft through meter tampering, privacy breaches of consumer data, and coordinated attacks on demand response systems (Ebtia et al., 2024).
@@
State Estimation. AC State Estimation is a crucial process in SG systems, designed to determine the most likely steady-state operation of the power system based on field measurements. It serves as a fundamental data processing tool responsible for supporting and increasing system visibility while filtering errors that may appear in real-time measurements, system topology, and parameters (Acurio et al., 2023).
@@
However, despite its significance, AC State Estimation is vulnerable to cyber threats, particularly FDIAs. These attacks can manipulate measurement data, introduce errors in state estimation, which can impact bus power flow calculations and the economic dispatch of the power system (Huang and Li, 2022). Within the FL framework, Lin et al. (Lin et al., 2022) propose an edge-based FL framework for detecting FDIAs in SG state estimation. This approach leverages local system operating data to predict and assess system stability in a distributed manner. By utilizing edge devices for local model training, the framework reduces the computational burden on central servers and enhances privacy protection. A key innovation in this work is the incorporation of an incentive mechanism designed to encourage high-quality data sharing and model training. This mechanism addresses the challenge of motivating participants to actively contribute to the FL process, which is essential for improving the overall detection accuracy of the system.
Events Classification and Anomaly Detection. Grid event classification and anomaly detection are vital for maintaining SG reliability. Deep learning models are often used but are known to be vulnerable to traditional adversarial attacks, where attackers make subtle changes to input data to cause misclassifications, a common concern even in centralized systems (Niazazari and Livani, 2020).
Addressing these issues, Mohammadabadi et al. (Mohammadabadi et al., 2023) investigated the impact of such adversarial attacks (e.g., using the Fast Gradient Sign Method (FGSM)) on their proposed federated XGBoost model for event classification with Phasor Measurement Unit (PMU) data. FGSM and similar techniques craft malicious inputs by adding small, calculated perturbations to clean data to deceive machine learning models. While their framework uses FL for privacy and homomorphic encryption to secure communications, their consideration of FGSM directly engages with how these established attack methods affect distributed learning systems. They also explored data processing techniques as a defense. Separately, Jithish et al. (Jithish et al., 2023) utilized FL for real-time anomaly detection in smart meters, focusing on privacy. To protect the integrity and confidentiality of model updates during transmission—a fundamental security concern in any distributed network—they employed standard secure communication protocols like SSL/TLS.
@@
Voltage Control. Voltage regulation is a critical process in electrical power systems, designed to maintain voltage levels within acceptable limits across the network. It involves controlling and adjusting the voltage to ensure stable and reliable power delivery to consumers (Xie et al., 2021). However, these networks are susceptible to data-related attacks, which can lead to voltage instability. Attackers may exploit vulnerabilities in communication channels to inject false data or disrupt the flow of information critical for effective voltage control.
@@
In the FL-based SG system, Liu et al. (Liu and Wu, 2022) combine FL with reinforcement learning techniques to develop a privacy-preserving and efficient control strategy. In this federated reinforcement learning framework, grid-connected interface inverters of renewable energy sources are modeled as deep neural network-based agents that interact with their local environment to learn optimal voltage-var control strategies. The process operates in a two-stage process: local training using reinforcement learning algorithms, followed by periodic global aggregation of model updates. This approach offers several advantages, including enhanced data privacy, reduced communication costs, improved scalability, and adaptability to changing network conditions.
4.1.3. Vulnerabilities in Consumption Stage
@@
The consumption part of an SG represents the end-user side of the energy distribution system (Kaur et al., 2020), encompassing homes, businesses, industries, and more. It integrates advanced technologies like AMI with smart meters for real-time usage monitoring and demand-side management to optimize energy consumption patterns. This system leverages IoT devices and real-time analytics to enhance energy efficiency and provide consumers with detailed insights into their usage. Additionally, sophisticated forecasting techniques and machine learning algorithms are employed to predict and manage energy loads effectively, ensuring a balance between supply and demand. However, smart meters and IoT devices, which are integral to these systems, can be vulnerable to attacks (Liu et al., 2021), potentially allowing malicious customers to manipulate energy consumption data or disrupt power supply.
@@
Load Forecasting. Smart meter measurements, while essential for accurate demand forecasting, raise significant privacy and data breach concerns for consumers. The fine-grained load data collected by smart meters can potentially reveal sensitive information about residential users’ daily routines and behaviors, posing risks to property security (Husnoo et al., 2023). In the FL framework, Husnoo et al. (Husnoo et al., 2023) propose a secure FL approach designed for residential short-term load forecasting. This approach addresses the critical balance between accurate demand prediction and protecting consumer privacy in smart meter data. FL allows collaborative model training without exposing raw data, as models are trained locally, and only parameters are shared with a central server. To enhance security, the framework incorporates advanced techniques like gradient quantization using SignSGD (Bernstein et al., 2018) and differential privacy (Dwork, 2006). Additionally, it employs hierarchical and peer-to-peer strategies to improve both privacy and accuracy. The system tackles challenges such as non-IID data distribution and potential attacks through methods like deep federated adaptation and secure aggregation. Experimental results demonstrate that this approach achieves high forecasting accuracy while preserving user data privacy, with significant improvements in prediction accuracy compared to traditional methods.
@@
Energy Theft Detection. Energy theft detection (Salinas and Li, 2015) in SGs refers to the process of identifying and preventing unauthorized consumption of electricity, an illegal activity that causes significant economic losses to utility companies. SGs integrate digital technologies with traditional power systems, enabling advanced energy management and monitoring capabilities. These systems utilize AMI and smart meters to collect, measure, and analyze energy usage data from customers at regular intervals. However, the collection and analysis of consumer energy data create vulnerabilities, potentially allowing attackers to compromise smart meters.
@@
Within the FL framework, Wen et al. (Wen et al., 2021) introduce FedDetect, a novel privacy-preserving FL framework designed for energy theft detection in SGs. FedDetect leverages FL to train machine learning models on distributed devices, aggregating local model updates instead of raw data to protect user privacy. Ashraf et al. (Ashraf et al., 2022) propose FedDP, a privacy-protecting theft detection scheme for SGs using FL. This framework addresses the challenge of utilizing consumer energy data for theft detection while maintaining privacy protection. Zafar et al. (Zafar et al., 2023) develop a FL-assisted hybrid deep learning model for electricity theft detection in SGs. Their model combines FL techniques with deep learning to enhance security and reliability of electricity theft detection. Li et al. (Li et al., 2024a) present FUSE, a novel electricity theft detection framework that integrates FL with U-shape split learning. FUSE aims to address the challenges of data privacy and security in SG systems while effectively detecting electricity theft.
4.2. SOTA FL-based Attack and Defense Methods
@@
4.2.1. SOTA FL Attack
@@
(1) Byzantine Attack in FL. Byzantine attacks in FL involve malicious clients sending incorrect or misleading model updates to the central server, aiming to disrupt the training process and degrade model performance (Fang et al., 2020). These malicious local model updates are typically intentionally modified by clients controlled by an attacker. We can categorize several classical and commonly discussed Byzantine attacks based on their approach to manipulating updates.
Naive Attacks. In their simplest form, naive attack methods aim to disrupt the convergence of the global model without sophisticated manipulation. One such approach is the Random Attack, where malicious clients send randomized updates drawn from a Gaussian distribution, $N(\mu,\sigma^{2}\textbf{I}_{d})$. The goal is to choose $\mu$ and $\sigma$ carefully so these malicious updates resemble benign ones in their distribution, making them harder to detect. Similarly, the Noise Attack involves malicious clients perturbing their local model updates by adding Gaussian noise from $N(\mu,\sigma^{2}\textbf{I}_{d})$. Here, the selection of $\mu$ and $\sigma$ is crucial: too little noise won’t affect convergence, while too much will make the malicious updates obviously different from benign ones. Another straightforward technique is the Sign-flip Attack, which, instead of altering the magnitude of updates, simply flips the element-wise sign of the model updates to disrupt the aggregation process.
Advanced Attacks. Beyond these straightforward, naive approaches, more sophisticated Byzantine attacks employ calculated strategies to subvert FL. The Min-Max/Min-Sum Attack (Shejwalkar and Houmansadr, 2021) is a two-step process: first, attackers perturb the average of benign updates to create a malicious update. Then, for Min-Max, this update is optimized to ensure its maximum Euclidean distance from any benign update doesn’t exceed the maximum distance between any two benign updates. In contrast, for Min-Sum, the malicious update is optimized so that the sum of its distances to each benign update is bounded by the maximum total distance of a benign update from others. These attacks are potent because malicious clients strategically search for optimal perturbations to maximize their impact.
Another specialized attack is the AGR-tailored Trimmed-mean Attack (Shejwalkar and Houmansadr, 2021). As proposed by Shejwalkar and Houmansadr (2021), this attack specifically targets defense mechanisms like Trimmed-mean (Yin et al., 2018a) by maximizing the Euclidean distance between the aggregated results of a simple average and the Trimmed-mean defense. While effective and intuitive, the malicious updates generated by this method significantly deviate from benign ones, making them relatively easy for advanced defense mechanisms to detect.
In contrast, the A-little-is-enough (Lie) Attack (Baruch et al., 2019) aims for stealth. Malicious clients apply only a slight perturbation to their local benign updates, making detection difficult. Specifically, they calculate the element-wise mean ($\mu_{j}$) and standard error ($\sigma_{j}$) of all updates, then generate elements of malicious updates as $(\beta_{i})_{j}=\mu_{j}-z\times\sigma_{j}$, where $\beta$ represents the malicious updates, $j$ is the element index, and $z$ is a scaling factor dependent on the proportion of malicious clients.
Finally, the ByzMean Attack (Xu et al., 2022) manipulates the mean of updates to an arbitrarily malicious value. It achieves this by dividing malicious clients into two groups. The first group generates malicious updates using any existing attack method (e.g., Lie attack). The second group then crafts their updates to ensure that the overall average of all client updates (benign and malicious) matches the average of the malicious updates from the first group.
(2) Backdoor Attack in FL. Backdoor attacks in FL aim to maintain the global model’s performance on clean inputs while causing it to make incorrect predictions on inputs containing a specific predefined feature (also called a trigger). Since backdoor attacks (Lyu et al., 2024, 2023) do not significantly affect the main task accuracy, the malicious local model updates are similar to benign ones, making anomaly detection more challenging in FL. Several classical and commonly discussed backdoor attacks are as follows.
Badnet Attack (Gu et al., 2017) is a simple and effective backdoor attack where the attacker injects the same trigger into the training data of compromised clients. The ground truth label of these training data will also be changed to another backdoor class. In contrast to Badnet, the Distributed Backdoor Attack (DBA) (Xie et al., 2019) sees the adversary decompose the backdoor trigger into several small triggers and apply them to different compromised clients. This distributed approach makes the attack harder to detect, as each client’s updates appear less suspicious. Building upon Badnet, the Scaling Attack (Bagdasaryan et al., 2020) involves malicious clients scaling their model updates in a way that amplifies the impact of their backdoor-triggered updates without affecting clean data performance. The scaling factor ensures the backdoor pattern has a stronger influence on the global model. The Projected Gradient Descent (PGD) Attack (Wang et al., 2020), adapted from adversarial machine learning, iteratively perturbs the model updates of compromised clients within a certain bound to embed the backdoor. This method is effective at creating small but impactful changes to the model parameters. Finally, the Neurotoxin Attack (Zhang et al., 2022) explores a persistent attack method designed for scenarios where attackers can only participate in a limited number of rounds. In this attack, malicious clients project their model updates into a subspace that excludes the global coordinates. This approach ensures that the projected updates from malicious clients are primarily embedded in coordinates with minimal perturbation from benign updates, maintaining the backdoor effect even after the attack ceases.
@@
(3) Inference Attack in FL. Membership Inference Attack (MIA) (Hayes et al., 2019; Shokri et al., 2017) is a major privacy attack that seeks to determine whether a specific data record was part of the training dataset used to build a machine learning model. In the context of FL, MIA poses a significant threat to the privacy of participating clients and their data. Various entities, including malicious clients, servers, or external attackers, can execute MIA in FL. These adversaries exploit the model’s natural tendency to overfit on training data, leading to unintended information leakage. Approaches to conducting MIA in FL include attack models (Nasr et al., 2019) and gradient-based MIAs, which exploit the gradients exchanged during the FL process. Passive MIA techniques also exist, where adversaries infer membership information by merely observing communication between clients and the server. Successful MIAs compromise the privacy of individual users, a particularly serious issue in sensitive domains such as healthcare or finance, where data confidentiality is paramount. Beyond identifying membership, MIA can also reveal attributes or properties of the training data, further threatening user privacy. Defending against MIA in FL presents unique challenges due to the distributed nature of the learning process, with traditional privacy-preserving techniques either being inapplicable or significantly impacting model utility. As a result, researchers are actively developing novel defense mechanisms to improve FL’s resilience to MIA while preserving model performance.
4.2.2. SOTA Defenses
To defend against the attacks mentioned above, the owner of the FL system can adopt various defense methods. In the following, we introduce classical and commonly discussed defense methods. To defend against Byzantine and backdoor attacks, the server can leverage defense methods, which can be generally divided into two main categories: detection-filtering methods (Blanchard et al., 2017; Bhagoji et al., 2019; Guerraoui et al., 2018; Xu et al., 2025b, 2022; Shejwalkar and Houmansadr, 2021; Cao et al., 2020; Xu et al., 2025d, c) and impact-reduction methods (Yin et al., 2018a; Fung et al., 2020; Huang et al., 2024; Pillutla et al., 2022; Panda et al., 2022; Zhang and Hu, 2023). To defend against inference attacks, one can primarily leverage differential privacy (DP) (Zhu et al., 2020; McMahan et al., 2017b; Hu et al., 2023a; Shi et al., 2023; Xu et al., 2025a).